Two regular quadrilateral vinyl tiles each of 1ft side overlap each other such that overlapping region is a regular octagon. What is the area of the overlapping region?
1
Expert's answer
2019-08-20T09:11:50-0400

Two regular quadrilaterals (squares) ABCD and EFGH each of 1 ft side: AB=BC=CD=DA=EF=FG=GH=HE=1 ft, angle A=B=C=D=E=F=G=H=90o. Squares ABCD and EFGH overlap each other such that overlapping region is a regular octagon IJKLMNPR (see Pic.).



The angle "\angle" AKL is the external angle for the angle "\angle" LKJ. Angle "\angle" ALK is the external angle for the angle "\angle" KLM. External angle + angle = 180 °. Since all angles (in particular, the angle "\angle" LKJ = "\angle" KLM="\angle" LMN=135 ° ) in the regular octagon are equal, then the external angles "\angle" AKL="\angle" ALK ="\angle" FLM= "\angle"FML= 45 ° are equal.

In the triangle △ ALK, two angles are equal, and the third angle is right angle. Triangle △ ALK is a isosceles right triangle.  In the triangle △ FLM, two angles are equal, and the third angle is right angle. Triangle △ FLM is a isosceles right triangle. All sides (in particular, KL=LM) in the regular octagon are equal, Triangles "\triangle" ALK and "\triangle" FLM are congruent by ASA theorem, hence AL=FL, AK=FM. The resulting triangles △ ALK, LFM, MBN, NGP, PCR, RHI, IDJ, JEK are isosceles right triangles: 

sides KA=AL=LF=FM=MB=BN=NG=GP=PC=CR=RH=HI=ID=DJ=JE=EK=leg.


The formula for finding the area of a regular octagon:

area of a regular octagon = 8 * base * height / 2.

OS is a height of the regular octagon. The height OS is half the side of the square ABCD:

OS=AB/2=1/2=0.5 ft.

NP is the base of a regular octagon. 

From the isosceles right triangle NGP, we express the base NP using the side NG (leg), applying the Pythagorean theorem:

NP is hypotenuse of triangle NGP;

leg2+leg2=hypotenuse2

2*leg2=hypotenuse2

"leg=\frac{1}{\sqrt{2}} *hypotenuse"

Express the side of square BC using the base of octagon NP and the sides of the triangles BN, PC (BN = PC = NG=leg):

BN+NP+PC=1 ft

leg+hypotenuse+leg=1

2*leg+hypotenuse=1.

Substitute the resulting expression obtained for leg into the previous equation:

"2*\frac{1}{\sqrt{2}} *hypotenuse+hypotenuse=1"

"hypotenuse*({\sqrt{2}}+1)=1"

"hypotenuse=\frac{1}{\sqrt{2}+1}=\frac{\sqrt{2}-1}{(\sqrt{2}-1)*(\sqrt{2}+1)}="

"=\frac{\sqrt{2}-1}{2-1}=\sqrt{2}-1"=0.414 ft is the base of a regular octagon.


The formula for finding the area of a regular octagon:


Area of a regular octagon = 8 * base * height / 2 = 8 * 0.414 * 0.5/2 =0.828 ft2.


Answer: the area of the overlapping region is 0.828 ft2.